Sharon is investing her money. She thinks that she should make $10 for every $150 she invests. How much she does expect to make on an investment of $5850?
Answer by Theo(13342)

10/150 = x/5850

solve for x to get x = 10*5850/150 = 390.

10/150 = .06666666..... = 1/15

1/15 * 5850 = 390

the ratio of 390/5850 is the same as the ratio of 10/150.

she expect to make 390.

if you divide 5850 by 150, you get 39

she gets 39 * 10 = 390.

all these different ways of looking at it yield the same result.